OpenStudy (anonymous):

Your ship leaves port on a bearing of N25E for 3.5 hours, at which time you change your course to N90E for 5.25 hours. If your constant speed is 20mph what bearing must you take to get back to port? How long will the return trip take?

OpenStudy (dumbcow):

 use Law of cosines to find length x use Law of Sines to find angle a

OpenStudy (anonymous):

Thank you so much for the diagram and help. I really appreciate it! Would my x=27.58 or did I miss calculate? Would a = 41.8 degrees?

OpenStudy (dumbcow):

no i am getting different answers \[x = \sqrt{70^2 + 105^2 -2(70)(105)\cos(115)}\] \[\sin a = \frac{70 \sin(115)}{x}\]

OpenStudy (anonymous):

Okay. So x should be 63.44 and how would you get a?

OpenStudy (dumbcow):

no not 63.44 , look at diagram, "x" is the longest side so it must be greater than 105

OpenStudy (anonymous):

My bad I miss read x= 148.78 then. and a should be the 63.44. Trig is really not my best subject

OpenStudy (dumbcow):

yes on x no on "a" sin(115) = .9063 sin(a) = .9063*(70/148.78) = .42639 a = 25.24

OpenStudy (anonymous):

Okay. Thank you so much! I sincerely appreciate all of your help. So this means my bearing for back to port would be 25.24 degrees SE? and it would take be 148.78/20 to return.

OpenStudy (dumbcow):

SW not SE not sure on bearing format W 25.24 S ?? or S 64.76 W

OpenStudy (anonymous):

Okay. Awesome! I will put down both. How did you come up with the 64.76? Thank you so much!

OpenStudy (dumbcow):

90 - 25 = 65

OpenStudy (anonymous):

Thank you!